an aero plane of mass 30000 kg travels in a horizontal loop of radius 10m at the rate of 1000 km/h. calculate the centripetal acceleration and centripetal force of aero plane
1
Expert's answer
2018-07-30T14:07:08-0400
a) We can find the centripetal acceleration from the formula:
a_c=v^2/r=(1000 km/h∙(1000 m)/(1 km)∙(1 h)/(3600 s))^2/(10 m)=7.716∙〖10〗^4 m/s^2 .
b) We can find the centripetal force from the formula:
F_c=ma_c=30000 kg∙7.716∙〖10〗^4 m/s^2 =2.31∙〖10〗^9 N.